The length of a rectangle is increased by 20 percent and its width is increased by fifty percent. By what percent does the area increase?
(A) 70 (B) 80 (C) 100 (D) 170 (E) 180

Answer by dabanfield(803)

The length of a rectangle is increased by 20 percent and its width is increased by fifty percent. By what percent does the area increase?
(A) 70 (B) 80 (C) 100 (D) 170 (E) 180

Let L be the original length and W the original width. The area of the rectangle is L*W.
By increasing the Length to 1.2*L and the width to 1.5*W the new area is:
(1.2)L*(1.5)W = (1.2*1.5)(L*W) = 1.8*(L*W)= 1.8*A = A + .80*A
So the answer is (B) 80